Imagine that there are currently 10,000 students enrolled at your institution. The school decides to increase tuition, and enrollment falls to 9,000. Tuition started at $4,000 per semester but has since gone up to $4,800. What is the elasticity of demand?


1
Expert's answer
2021-03-17T19:36:34-0400

"Q_1 = 10000 \\

Q_2 = 9000 \\

P_1 = 4000 \\

P_2 = 4800"

Price elasticity of demand (using mid point formula):

"E = \frac{Q_2 – Q_1}{\frac{1}{2}(Q_2+Q_1)} \div \frac{P_2-P_1}{\frac{1}{2}(P_2+P_1)} \\

= \frac{9000 – 10000}{\frac{1}{2}(9000+10000)} \div \frac{4800-4000}{\frac{1}{2}(4800+4000)} \\

= \frac{-1000}{9500} \div \frac{800}{4400} \\

= -0.578"